OpenStudy (anonymous):

solve the initial value problem by the Laplace transform method: y''+4y=0 ; y(0)=1, y'(0)=1 solve the initial value problem by the Laplace transform method: y''+4y=0 ; y(0)=1, y'(0)=1 @Mathematics

OpenStudy (nikvist):

\[y''+4y=0\quad;\quad y(0)=y'(0)=1\]\[s^2Y-sy(0)-y'(0)+4Y=0\]\[(s^2+4)Y=sy(0)+y'(0)=s+1\]\[Y=\frac{s+1}{s^2+4}=\frac{s}{s^2+4}+\frac{1}{s^2+4}=\frac{s}{s^2+2^2}+\frac{1}{2}\cdot\frac{2}{s^2+2^2}\]\[y=\cos{2t}+\frac{\sin{2t}}{2}\]

OpenStudy (anonymous):

is there a way to check these through wolfram that you know of?

OpenStudy (anonymous):

not your answer....my other work